Benzaldehyde is a colorless liquid with the sweet smell of almond extract. A 50-mL sample of benzaldehyde weighs 52 grams. What is a density of benzaldehyde (include units in answer)? Is it more dense or less dense than water?
1
Expert's answer
2020-01-27T05:15:56-0500

A density is calculated by the following formula:

d = m / V,

where d - density, m - mass, V - volume.

A density of benzaldehyde is:

d = 52 g / 50 ml = 1.04 g/ml = 1,040 kg/m3

As the maximal water density equals 1,000 kg/m3, benzaldehyde is more dense than water.


Answer: 1,040 kg/m3